OpenStudy (anonymous):

Solve using the quadratic formula. 4z2 + 9z + 4 = 0

OpenStudy (anonymous):

\[x = \frac{-b \pm \sqrt{b-4ac}}{2a}, a = 4, b = 9, c = 4\]

OpenStudy (anonymous):

z = [-9 +- sqrt(81 - 4*4*4)] / 8 = [-9 +- sqrt(17) ] / 8

OpenStudy (anonymous):

(-9+sqrt(17))/8 and (-9-sqrt(17))/8

OpenStudy (anonymous):

x = -0.609611... or -1.6403...

OpenStudy (anonymous):

thanks guys